# A Community Palm Model

### v1

@@ Abstract
Palm oil production has been identified as one of the major drivers of deforestation for tropical countries. To meet supply chain objectives, soft commodity producing companies and other stakeholders need timely information of land cover dynamics in their supply shed. However, such data are difficult to obtain from suppliers who may lack digital geographic representations of their supply sheds and production locations. There is also a proliferation of mapping products coming onto the market, which have a spectrum of methods, definitions and geographic extents that may present conflicting information and can quickly become outdated. Here we present a “community model,” a machine learning model trained on pooled data sourced from many different stakeholders, to develop a specific land cover probability map, in this case a semi-global oil palm map. An advantage of this method is the inclusion of varied inputs, the ability to easily update the model as new training data becomes available and run the model on any year that input imagery is available. Inclusion of diverse data sources into one probability map can help establish a shared understanding across stakeholders on the presence and absence of a land cover or commodity (in this case oil palm). The model predictors are annual composites built from publicly available satellite imagery provided by Sentinel-1, Sentinel-2, and ALOS DSM. We provide map outputs as the probability of palm in a given pixel, to reflect the uncertainty of the underlying state (palm or not palm). The initial version of this model provides global accuracy estimated to be approximately 90% (at 0.5 probability threshold) from spatially partitioned test data. This model, and resulting oil palm probability map products are useful for accurately identifying the geographic footprint of palm cultivation. Used in conjunction with timely deforestation information, this palm model is useful for understanding the risk of continued oil palm plantation expansion in sensitive forest areas.
@@ 1 Introduction
Commodity supply chains have come to global attention in recent years, driven by heightened awareness of deforestation and its consequential CO2 emissions resulting from land cover conversion (Pendrill et al., 2019, 2022). This revelation has spurred voluntary commitments from major consumer packaged goods suppliers, exemplified by Unilever’s engagement, which has pledged to pursue deforestation-free practices (Vijay et al., 2016). Initiatives like the Round-table on Sustainable Palm Oil (RSPO) serve as proactive measures undertaken by industry stakeholders to address the environmental challenges associated with palm oil production. The urgency for accurate and timely information about commodity production locations has been further emphasized by recent regulations such as the European Union Regulation on Deforestation-free products (EUDR). Notably, the European Union identified palm oil as the primary driver of Union-driven deforestation among the eight commodities analyzed (European Commission, 2023).
Regulations such as the European Union Deforestation Regulation (EUDR), the United Kingdom’s Environment Act, and proposed bills like the US FOREST Act, are pushing deforestation-free commodity production from voluntary commitments to a regulatory requirement for doing business in global markets. With these regulations, traceability of commodities to the location of production, and due diligence demonstrating deforestation free sourcing backed by verifiable data, is required. In the case of the EUDR, this means organizations operating or placing products in the EU market must disclose sourcing locations and ensure full traceability of palm oil (and other commodities) as part of their due diligence process (EUDR Article 4, European Commission 2023). To meet EUDR due diligence requirements operators must collect information about commodity production, including the geolocation, and perform a risk assessment that demonstrates that the commodities are deforestation free (EUDR Article 8-10, European Commission 2023). This necessitates large-scale, wall-to-wall, robust, and regularly updated information on land cover and land use (European Commission, 2023).
Remote sensing and machine learning techniques offer promising methodologies for rapidly creating large scale land cover commodities datasets to support organizations meeting deforestation due diligence in their supply chains. Although commodity probability maps are not solely sufficient to ensure compliance, they are necessary to support organizations’ understandings of the impact of their supply chains, verify information provided by producers, and support due diligence reporting.
@@
Remote sensing-based mapping of oil palm is an area of active research. A combination of optical and synthetic aperture radar (SAR) data is often employed with a classifier, such as random forest, to outperform either data source alone (Nomura and Mitchard, 2018; Ordway et al., 2019; Descals et al., 2019; Sarzynski et al., 2020; Xu et al., 2020; Abramowitz et al., 2023). Li et al. (2017) used auto-encoders on high resolution QuickBird imagery to detect and count individual oil palm trees in Malaysia. Working with Sentinel-1 and Sentinel-2 data, Descals et al. (2021) used a convolutional neural network (CNN) to map oil palm globally, with distinction between industrial and smallholder plantations. Danylo et al. (2021) applied unsupervised methods to Landsat and Sentinel-1 data to map oil palm over time. In conjunction with maps constructed from remotely sensed imagery, Gaveau et al. (2018) and Gaveau et al. (2022) used manual delineation by experts to map oil palm.
The aim of this study is to leverage existing work and diverse data sources to develop a community supported Machine Learning (ML) model capable of generating annual oil palm maps from the most recent data available. Building on the existing studies mentioned above, we used the published data (where available) from these and other sources to create a communal training/validation dataset and a community ML model – a model that can be employed on-demand to create oil palm maps for time periods and locations with suitable input data. The advantage of this type of model is that it leverages existing work, is inclusive of multiple datasets from multiple contributors, improves over time, leverages public imagery sources and doesn’t lead to a proliferation of competing products. We characterize the accuracy of the community palm model, and show that estimated probabilities output by the model can be used in statistically-based decision making frameworks. Specifically, we use the model output at two points in time in a risk analysis framework, to quantify the risk of palm transitions in forested areas. This risk assessment is designed to elucidate the probability of palm conversion either from, or to, other land covers in the time and area of interest. We make the model, input data and output data available to the community for ongoing collaborative development.
@@ 2 Methods
The methods were chosen to help organizations understand the location of oil palm production, and change in such production over time as it relates to deforestation. The workflow is designed to be iterative, to facilitate rapid acquisition of additional training/validation data, model retraining and model deployment. To build this workflow, reference data were ingested into Google Earth Engine from multiple sources, and then overlaid on geographic predictors (imagery) to create training and validation sets. A TensorFlow-based neural network (Abadi et al., 2015) was trained on the data and deployed to Google Vertex AI for inference at scale in Earth Engine. The model accuracy was assessed with the validation data. We performed a cross-source validation, which retrained the model leaving out one training dataset at a time to understand the contribution of each training set to the overall model performance. We also present a method to estimate risk of palm-driven deforestation by looking at predictions at two points in time for the same area.
@@ 2.1 Training/Testing data
A key to this approach is the ongoing community contribution of the training and testing data that we used to build the model. Initial data sources were provided by community contributors to the Forest Data Partnership. Accuracy improves over time as additional community members contribute data to the model, and it can be re-trained and republished on demand. The data described below are provisional, and we expect stakeholders to provide more data in their areas of interest, with commensurate increase in model accuracy in those areas. Additional data sources will be cited as new model versions are published.
@@
Using Maxar WorldView-3 for visual reference, Google collected 8054 1024x1024 patches at 0.3 meter resolution at nadir, independently hand-labeled by two separate annotators. Oil Palm was defined as “oil palm trees planted in rows or terraces for agricultural purposes.” Bare ground patches of yet-to-be-planted palm plantations were not labeled as palm. These were sampled to generate 374,984 points at 10 meter resolution of palm frequency in [0, 1] based on the proportion of the 10 meter pixel annotated as palm. These annotations were performed on recent imagery available at the time of annotation, and were assumed to fairly represent 2021 conditions. Palm plantations operate on a growth and replanting cycle of approximately 25 years, leading to the assumption that annotation of imagery from a few years prior to 2021 would generally reflect the state of palm in 2021.
@@
Vollrath et al. (2019) provided 4,126 training and 1019 testing patches as 120x120 meter polygons representing palm in 2017. These patches were sampled and pooled to produce 296,049 points at 10 meter resolution: palm (1) or not palm (0).
@@
Publicly available validation data from Danylo et al. (2021) were downloaded from their GitHub site in the form of 10,303 palm presence (1) or absence (0) points (in 30x30 meter patches). These resulted in 7624 10-meter samples of palm (1) or not palm (0) in 2017. Publicly available validation data from Descals et al. (2021) were downloaded from their Zenodo site in the form of 13,495 palm presence (1) or absence (0) points (in 10x10 meter patches). These resulted in 11,531 10-meter samples of palm (1) or not palm (0) in 2019. Points for which the the input imagery was masked in any band were discarded, resulting in the difference in number between reference and resultant training points.
@@
T. Lips (personal communication) provided 180 labeled points according to palm presence or absence in 2023. These were hand-digitized using an earlier version of the model output as reference. Specifically, points of palm presence were placed in areas of obvious underprediction (false negatives) and points of palm absence were placed in areas of obvious overprediction (false positives).
@@ 2.2 Predictor data
The predictors are annual composites built from publicly available satellite imagery provided by Sentinel-1, Sentinel-2 and ALOS DSM. Specifically, the following Sentinel-2 Top-of-atmosphere (TOA) reflectance bands were used in order to present a high density of information in areas of spectral importance: B1, B2, B3, B4, B5, B6, B7, B8, B8A, B9, B10, B11, B12. TOA reflectance data were used due to the increased availability of TOA data compared to data products with a higher level of processing, surface reflectance for example. We built annual composites by taking the cloud-masked pixel- and band-wise mean for all available imagery in a calendar year. The Cloud Score+ dataset (Pasquarella et al., 2023) was used for the cloud mask, with a 0.6 threshold on the ’cs_cdf’ band. The means were scaled to reflectance in approximately [0, 1] and no further normalization was performed.
Sentinel-1 data were selected from interferometric wide-swath scenes that included both VV (vertical-vertical) and VH (vertical-horizontal) polarizations in both ascending and descending orbital paths. Radiometric compensation for terrain using the volume scattering model described by (Vollrath et al., 2020) was used to preprocess backscatter and the following statistics computed for a calendar year: minimum, maximum, mean, and standard deviation. The statistics were stretched to decibels and subsequently normalized at train time as described below.
@@
The eight Sentinel-1 statistics (min, max, mean, and SD for both VV and VH) were stacked with slope and the 13 Sentinel-2 bands in a given calendar year to create the input composites.
The palm datasets described above were combined with composites from the relevant years to produce 979,585 multi-temporal training/testing points spanning 41 countries and eight biomes (Dinerstein et al., 2017), with class distribution:
  • 435,077 negatives (not palm)
  • 46,811 maybes (partial palm annotation in 10 m pixel, where 0% < palm % < 100%)
  • 611,666 positives (palm)
The geographic distribution of reference data is illustrated in Figure 1:
Figure 1: Location of training/testing data points (white).
@@ 2.3 Model training and inference
The model is a simple neural network with a single hidden layer of 64 nodes, a dropout layer (at 10%) and a sigmoid output. This model was chosen for its simplicity in terms of number of parameters, but also flexibility in functional representation given a vector input (i.e. a per-pixel model in which the input is 1x1xC for C bands). The training data are treated as a frequency distribution and compared directly to the model output using binary cross-entropy loss. The statistics of the training set were computed and used to standardize Sentinel-1 data coming into the model to have unit mean and variance (with a normalization layer). The Sentinel-2 data were input as TOA reflectance in [0, 1]) and slope as normalized degrees in [0, 1]. The model was fit to the training data over seven epochs. No hyperparameter tuning, ablation analysis or other optimizations were performed. The output of the model (F(xt)) is the estimated conditional probability of palm presence given the vector of covariates (xt) at time t.
@@ 2.4 Model accuracy assessment
The accuracy of the model was estimated through a cross-validation process over several partitions of the training data. The pooled data were geographically partitioned into three folds based on a hexagonal decomposition of an equal-area projection, where each cell was roughly 260 square kilometers. See Goldblatt et al. (2018) and Fairfax et al. (2023) for a detailed description of this decomposition. Because the total boundary length of a hexagonal partition is minimized, we chose this method to minimize spatial autocorrelation between training and validation data. The distribution of folds (grayscale) and reference data locations (yellow) over a portion of Malaysia are illustrated in Figure 2.
Figure 2: Distribution of folds (grayscale) and reference data locations (yellow).
We used three approximately equal geographic folds (n $\approx$ 305,151) in a three-fold cross-validation (Efron and Gong, 1983). Each of the three folds was withheld as testing and the remaining two folds used as training. The following accuracy measures are reported for each of the three folds: binary cross-entropy loss, binary overall accuracy at 0.5 threshold of output, area under the receiver operator characteristics (ROC) curve (AUC) (Fawcett, 2006), precision at 0.95 recall, recall at 0.95 precision, recall and precision at 0.5, 0.7 and 0.9 thresholds on the output. All measures were computed according to the Keras metrics implementations in TensorFlow.
The training data were also partitioned by the dataset provider in a leave-one-group-out cross-validation. As noted in (Zhao et al., 2014), "self-evaluation" is a notorious overestimate of map accuracy. This partition was intended to fairly assess the self-evaluation effect. In this partition, the two training folds (from the geographic partition) of all providers except one were pooled, a model trained, and its output compared to both the training and validation folds (geographic partition) of the left-out provider.

### v2

@@ Abstract
Palm oil production has been identified as one of the major drivers of deforestation for tropical countries. To meet supply chain objectives, soft commodity producing companies and other stakeholders need timely information of land cover dynamics in their supply shed. However, such data are difficult to obtain from suppliers who may lack digital geographic representations of their supply sheds and production locations. There is also a proliferation of mapping products coming onto the market, which have a spectrum of methods, definitions and geographic extents that may present conflicting information and can quickly become outdated. Here we present a “community model,” a machine learning model trained on pooled data sourced from many different stakeholders, to produce a map of palm probability at global scale. An advantage of this method is the inclusion of varied inputs, the ability to easily update the model as new training data becomes available and run the model on any year for which input imagery is available. Inclusion of diverse data sources into one probability map can help establish a shared understanding across stakeholders on the presence and absence of a land cover or commodity (in this case oil palm). The model predictors are annual composites built from publicly available satellite imagery provided by Sentinel-1, Sentinel-2, and ALOS-2, and terrain data from Jaxa (AW3D30) and Copernicus (GLO-30). We provide map outputs as the probability of palm in a given pixel, to reflect the model certainty of the underlying state (palm or not palm). This version of this model provides global accuracy estimated to be  92% (at 0.5 probability threshold) on an independent test set. This model, and resulting oil palm probability map products are useful for accurately identifying the geographic footprint of palm cultivation. Used in conjunction with timely deforestation information, this palm model is useful for understanding the risk of continued oil palm plantation expansion in sensitive forest areas.
@@ 1 Introduction
Commodity supply chains have come to global attention in recent years, driven by heightened awareness of deforestation and its consequential CO2 emissions resulting from land cover conversion (Pendrill et al., 2019, 2022). This revelation has spurred voluntary commitments from major consumer packaged goods suppliers, exemplified by Unilever, which has pledged to pursue deforestation-free practices (Vijay et al., 2016). Initiatives like the Round-table on Sustainable Palm Oil (RSPO) serve as proactive measures undertaken by industry stakeholders to address the environmental challenges associated with palm oil production. The urgency for accurate and timely information about commodity production locations has been further emphasized by recent regulations such as the European Union Deforestation Regulation (EUDR). Notably, the European Union identified palm oil as the primary driver of Union-driven deforestation among the eight commodities analyzed (European Commission, 2023).
EUDR, the United Kingdom’s Environment Act, and proposed bills like the US FOREST Act, are pushing deforestation-free commodity production from voluntary commitments to a regulatory requirement for doing business in global markets. In the case of the EUDR, this means organizations operating or placing products in the EU market must provide (European Commission, 2023):
  • Geo-referenced location(s) of commodity production (EUDR Article 9, European Commission 2023).
  • A due diligence process that demonstrates that the locations from which commodities are sourced are at low risk of deforestation (EUDR Article 10, European Commission 2023).
The risk assessment portion of the due diligence process requires large-scale, wall-to-wall, robust, and regularly updated information on land cover and land use in order to ensure that deforestation is not occurring in the sourcing locations. Remote sensing and machine learning techniques offer promising methodologies for rapidly creating large scale land cover commodity datasets to support organizations’ due diligence reporting on their supply chains. Although commodity probability maps are not solely sufficient to ensure compliance, they are necessary to support organizations’ understandings of the impact of their supply chains, verify information provided by producers, and support due diligence reporting.
@@
Remote sensing-based mapping of oil palm is an area of active research. A combination of optical and synthetic aperture radar (SAR) data is often employed with a classifier, such as random forest, to outperform either data source alone (Nomura and Mitchard, 2018; Ordway et al., 2019; Descals et al., 2019; Sarzynski et al., 2020; Xu et al., 2020; Abramowitz et al., 2023). Li et al. (2017) used auto-encoders on high resolution QuickBird imagery to detect and count individual oil palm trees in Malaysia. Working with Sentinel-1 and Sentinel-2 data, (Descals et al., 2021, 2024) used a convolutional neural network (CNN) to map oil palm globally, with distinction between industrial and smallholder plantations. Danylo et al. (2021) applied unsupervised methods to Landsat and Sentinel-1 data to map oil palm over time. In conjunction with maps constructed from remotely sensed imagery, Gaveau et al. (2018) and Gaveau et al. (2022) used manual delineation by experts to map oil palm.
The aim of this study is to leverage existing work and diverse data sources to develop a community supported Machine Learning (ML) model capable of generating annual oil palm maps from the most recent data available. Building on the existing studies mentioned above, we used the published data (where available) from these and other sources to create a communal training/validation dataset and a community ML model – a model that can be employed on-demand to create oil palm maps for time periods and locations with suitable input data. The advantage of this type of model is that it leverages existing work, is inclusive of multiple datasets from multiple contributors, improves over time, leverages public imagery sources, doesn’t lead to a proliferation of competing products and gives consumers access to the means of map production. We characterize the accuracy of the community palm model, and show that estimated probabilities output by the model can be used in statistically-based decision making frameworks. Specifically, we use the model output at two points in time in a risk analysis framework, to quantify the risk of palm transitions in forested areas. This risk assessment is designed to elucidate the probability of palm conversion either from, or to, other land covers in the time and area of interest. We make the model, input data and output data available to the community for ongoing collaborative development.
@@ 2 Methods
The methods were chosen to help organizations understand the location of oil palm production, and change in such production over time as it relates to deforestation. The workflow is designed to be iterative, to facilitate rapid acquisition of additional training/validation data, model retraining and model deployment. To build this workflow, reference data were ingested into Google Earth Engine (Gorelick et al., 2017) from multiple sources, and then overlaid on geographic predictors (imagery) to create training and validation sets. A TensorFlow-based neural network (Abadi et al., 2015) was trained on the data and deployed to Google Vertex AI for inference at scale in Earth Engine. The model accuracy was assessed with the validation data. We performed a cross-source validation, which retrained the model leaving out one training dataset at a time to understand the contribution of each training set to the overall model performance. We also present a method to estimate risk of palm-driven deforestation by looking at predictions at two points in time for the same area.
@@ 2.1 Training/Testing data
A key to this approach is the ongoing community contribution of the training and testing data that we used to build the model. Initial data sources were provided by community contributors to the Forest Data Partnership. Accuracy improves over time as additional community members contribute data to the model, which can be easily republished and re-evaluated. The data described below are provisional, and we expect stakeholders to provide more data in their areas of interest, with commensurate increase in model accuracy in those areas. Additional data sources will be cited as new model versions are published.
@@
Using Maxar WorldView-3 for visual reference, Google collected 8054 1024x1024 patches at 0.3 meter resolution at nadir, independently hand-labeled by two separate annotators. Oil Palm was described to annotators as “oil palm trees planted in rows or terraces for agricultural purposes.” Bare ground patches of yet-to-be-planted palm plantations were not labeled as palm.
These were sampled to generate 326,714 points at 10 meter resolution of palm frequency in [0, 1] based on the proportion of the 10 meter pixel annotated as palm or not-palm by both annotators. These annotations were performed on recent imagery available at the time of annotation, and were assumed to fairly represent 2021 conditions. Palm plantations operate on a growth and replanting cycle of approximately 25 years, leading to the assumption that annotation of imagery from a few years prior to 2021 would generally reflect the state of palm in 2021.
@@
Vollrath et al. (2019) provided 5,145 patches as 120x120 meter polygons representing palm in 2017. These patches were sampled and pooled to produce 738,421 points at 10 meter resolution: palm (1) or not palm (0).
@@
Publicly available validation data from Danylo et al. (2021) were downloaded from their GitHub site in the form of 10,303 palm presence (1) or absence (0) points (in 30x30 meter patches). These resulted in 10,265 10-meter samples of palm (1) or not palm (0) in 2017.
@@
T. Lips (personal communication) provided 180 labeled points according to palm presence or absence in 2023. These were hand-digitized using an earlier version of the model output as reference. Specifically, points of palm presence were placed in areas of obvious under-prediction (false negatives) and points of palm absence were placed in areas of obvious over-prediction (false positives).
Publicly available data for palm plantations in Peru for 2019-2020 were obtained from Fricker et al. (2022) in the form of 1288 polygons. These data were combined with tree cover masks, a stable forest mask, the Descals et al. (2021) palm mask and subsampling (0.005 rate) at 10 meter resolution to produce 6469 palm (1) presence points.
Publicly available validation data from Descals et al. (2024) were downloaded from their Zenodo site (https://zenodo.org/records/13379129) in the form of 18,812 palm presence (1) or absence (0) points (in 10x10 meter patches). These resulted in 18,736 10-meter samples of palm (1) or not palm (0) in 2021. These data were not used for training or model selection.
To balance the training data, pseudo-absence points (assumed not palm locations) were generated from ancillary layers on forest cover and land cover. Specifically, Dynamic World (Brown et al., 2022) was used to find stable non-tree areas, defined as majority non-tree classes over a three year period (2019-2021). Multiple forest datasets were combined to estimate stable, non-plantation, forest cover over the observation period (1984-2020). Palm pseudo-absences were generated by stratified sampling of stable non-tree and stable forest areas in the following regions: Southeast Asia (n=14,124), Central America (n=9624), West Africa (n=35,145), and Peru (n=7819).
@@ 2.2 Predictor data
The predictors are annual composites built from publicly available satellite imagery provided by Sentinel-1, Sentinel-2, and ALOS-2, and terrain data from Jaxa (AW3D30) and Copernicus (GLO-30). Specifically, the following Sentinel-2 Top-of-atmosphere (TOA) reflectance bands were used: B1, B2, B3, B4, B5, B6, B7, B8, B8A, B9, B10, B11, B12. TOA reflectance data were used due to the increased availability of TOA data compared to data products with a higher level of processing, surface reflectance for example. We built annual composites by taking the cloud-masked pixel- and band-wise mean for all available imagery in a calendar year. The Cloud Score+ dataset (Pasquarella et al., 2023) was used for the cloud mask, with a 0.6 threshold on the ’cs_cdf’ band. The means were scaled to reflectance in approximately [0, 1] and no further normalization was performed.
Sentinel-1 data were selected from interferometric wide-swath scenes that included both VV (vertical-vertical) and VH (vertical-horizontal) polarizations in both ascending and descending orbital paths. Radiometric compensation for terrain using the volume scattering model described by Vollrath et al. (2020) was used to preprocess backscattering coefficients. For a calendar year, minimum, maximum, mean, and standard deviation of backscatter were computed at each polarization and subsequently stretched to decibels.
Palsar-2 annual composites, containing HH and HV polarizations at 25 meters resolution (Shimada et al., 2014), were bilinearly resampled, gap filled with the rolling 3-year mean of annual composites, and converted to scaled decibels.
@@
In a given calendar year, the eight Sentinel-1 statistics (min, max, mean, and SD for both VV and VH) were stacked with the 13 Sentinel-2 band means, the two Palsar-2 bands (HH and HV) and slope to create the annual input composites.
The palm datasets described above were combined with composites from the relevant years to produce 979,585 multi-temporal training/testing points spanning 59 countries and 10 biomes (Dinerstein et al., 2017), with class distribution:
  • 575,893 negatives (not palm)
  • 532,467 positives (palm)
The pooled data were geographically partitioned into three folds based on a hexagonal decomposition of an equal-area projection, where each cell was roughly 26,000 square kilometers. See Goldblatt et al. (2018) and Fairfax et al. (2023) for a detailed description of this decomposition. Because the total boundary length of a hexagonal partition is minimized, we chose this method to minimize spatial autocorrelation between training, testing validation data. The size and label balance of the folds vary slightly, a property we leveraged by choosing the largest fold for training. Of the remaining two folds, one was used for model selection and hyperparameter tuning (validation) and the other for evaluation (i.e. not used in model training or selection).
The geographic distribution of positive (palm=1) and negative (palm=0)reference points is illustrated in Figure 1. The distribution of over a portion of Southeast Asia are illustrated in Figure 2.
Figure 1: Location of training/testing reference locations: negatives (white), positives (gray).
Figure 2: Distribution of folds: fold 0 (red), fold 1 (green), fold 2 (blue)
@@ 2.3 Model training and inference
The model is a neural network with four hidden layers and a single sigmoid output. It is a per-pixel model in which the input is 1x1xC (for C bands) vector and the output (F(xt)) is the estimated conditional probability of palm presence given the vector of covariates (xt) at time t. The architecture, activation function, optimization function, learning rate, batch size, training iterations and other parameters of the model and training configuration were chosen automatically using a hyperparameter search performed on infrastructure similar to Google Cloud AutoML (https://cloud.google.com/automl).
@@ 2.4 Model accuracy assessment
We estimate overall map accuracy in two ways. We used the test fold (not seen in training or model selection/tuning) to compute the following accuracy measures: binary cross-entropy loss, binary overall accuracy at 0.5 threshold of output, area under the receiver operator characteristics (ROC) curve (AUC) (Fawcett, 2006), recall and precision at 0.5 threshold on the output. As noted in (Zhao et al., 2014), "self-evaluation" is a notorious overestimate of map accuracy. We also use Descals et al. (2024) validation data to compute the same measures (all folds and only folds 0 and 1, which are not in the same geographic fold as the training data). The Descals et al. (2024) training or validation data were not used in model training or selection.